# On twisted Kawamata’s semi-positivity and finite generation of generalized canonical rings

### v1

@@ 1. Introduction
[…]
@@
We collect the notion of generalized pairs by Birkar in Section 5. When $K_{X}+B+M$ is big, the above theorem directly follows from [BCHM] since we perturb $B+M$ with $K_{X}+B+M$ and produce a usual klt boundary $B^{\prime}$ such that $B^{\prime}\sim_{\mathbb{Q}}B+M+\varepsilon(K_{X}+B+M)$ for a small $\varepsilon>0$. Thus, we apply the Fujino–Mori argument [FM]. Finally, we note the interesting difference between the usual and generalized pair in the Kawamata semi-positivity, the canonical bundle formula, and the finite generation. For generalized pairs, we do now know the functorial behavior of the moduli part of [Am04, Theorem 0.2 (1)]. For the case of (generalized) log Calabi–Yau fibrations, it is known by [Fil]. However, in general, it is unknown. Fortunately, this does not matter for our application. Moreover, one may ask the case of log canonical in Theorem 1.2. However, the finite generation of lc generalized canonical rings does not hold by [G, Remark 3.5].
@@
The first author was partially supported by grants JSPS KAKENHI ##\##18H01108, 19KK0345, 20H00111, 21H00970, 22H01112, 23H01064. The second author was partially supported by grants JSPS KAENHI ##\##21H00979. The first author thanks Professor Paolo Cascini for the discussion.
@@ 4. Semi-positivity of the moduli part
(2) The ${\mathbb{Q}}$-line bundle ${f^{\prime}}^{*}L$ admits a singular Hermitian metric $-~\widetilde{h}$ with semi-positive curvature, which is canonically constructed via the isomorphism ${f^{\prime}}^{*}L\sim_{{\mathbb{Q}}}K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}-\ulcorner{D^{\prime}_{-}}\urcorner$ and $b^{-1}$. The corresponding singular Hermitian metric $h_{Y}$ on $L$ (such that $-~{f^{\prime}}^{*}h_{Y}=\widetilde{h}$) has semi-positive curvature.
@@ 5. On finite generation problem of log canonical rings for generalized pairs
If $\kappa(X,l(K_{X}+B+M))\leq 0$, the finite generation is clear. Thus we may assume that $\kappa(X,l(K_{X}+B+M))>0$. We consider the Iitaka fibration $f:X\dashrightarrow Y$of $l(K_{X}+B+M)$ and take a higher birational model $X^{\prime}\to X$ induces a holomorphic morphism $f^{\prime}:X^{\prime}\to Y$ which is birational equivalent to $f$. Thus, by replacing, we may assume that $(X,B)$ is a simple normal crossing klt pair admitting the Iitaka fibration $f:X\to Y$ of $K_{X}+B+M$, and that $M$ has a semi-positive singular Hermitian metric with a vanishing Lelong number for the potential at every point of $X$. Now, we apply the weak semi-stable reduction theorem for $f$.
@@
We use the notation in twisted Kawamata’s semi-positivity theorem 1.1. Consider the divisor $\Delta_{Y^{\prime}}^{-}$. Since $B$ is effective, any components of $f^{\prime*}\Delta_{Y^{\prime}}^{-}$ are $\varphi$-exceptional divisors. Thus, $f^{\prime}$ is also the Iitaka fibarion of $K_{X^{\prime}}+f^{\prime*}\Delta_{Y^{\prime}}^{-}+B^{\prime}+M^{\prime}$. Therefore the finite generation of $R(K_{X}+B+M)$ is reduced to that of $R(K_{X^{\prime}}+f^{\prime*}\Delta_{Y^{\prime}}^{-}+B^{\prime}+M^{\prime}).$ And that is equivalent to the finite generation of $R(K_{Y^{\prime}}+M_{Y^{\prime}}+\Delta^{+}_{Y^{\prime}})$. Now $(Y^{\prime},\Delta_{Y^{\prime}}^{+})$ is klt and $M_{Y^{\prime}}$ is nef by Theorem 1.1. Siince $K_{Y^{\prime}}+M_{Y^{\prime}}+\Delta^{+}_{Y^{\prime}}$ is big, the desired finite generation follows from [BCHM]. ∎
The following is an interesting corollary for even smooth projective variety.

### v2

@@ 1. Introduction
$f^{*}Q_{l}$ $=$ $\sum_{j}w_{lj}P_{j},\ \text{where}\ w_{lj}\in{\mathbb{Z}}_{>0},$
$\bar{d}_{j}$ $=$ $\frac{d_{j}+w_{lj}-1}{w_{lj}}(\in{\mathbb{Q}}_{>1})\ \text{if}\ f(P_{j})=Q_{l},$
@@
We collect the notion of generalized pairs by Birkar in Section 5. When $K_{X}+B+M$ is big, the above theorem directly follows from [BCHM] since we perturb $B+M$ with $K_{X}+B+M$ and produce a usual klt boundary $B^{\prime}$ such that $B^{\prime}\sim_{\mathbb{Q}}B+M+\varepsilon(K_{X}+B+M)$ for a small $\varepsilon>0$. Thus, we apply the Fujino–Mori argument [FM]. Finally, we note the interesting difference between the usual and generalized pair in the Kawamata semi-positivity, the canonical bundle formula, and the finite generation. For generalized pairs, we do now know the functorial behavior of the moduli part of [Am04, Theorem 0.2 (1)]. For the case of (generalized) log Calabi–Yau fibrations, it is known by [Fil]. However, in general, it is unknown. Fortunately, this does not matter for our application. Moreover, one may ask the case of log canonical in Theorem 1.2. However, the finite generation of lc generalized canonical rings does not hold by [G, Remark 3.5]. After putting this paper on the arXiv, we know that Hacon and Păun are working with the similar idea with us in [HP]. They also treat various applications for the adjunction theory.
@@
The first author was partially supported by grants JSPS KAKENHI ##\##18H01108, 19KK0345, 20H00111, 21H00970, 22H01112, 23H01064. The second author was partially supported by grants JSPS KAENHI ##\##21H00979. The first author thanks Professor Paolo Cascini and Chi-Kang Chang for the discussion.
@@ 4. Semi-positivity of the moduli part
(2) The ${\mathbb{Q}}$-line bundle ${f^{\prime}}^{*}L$ admits a singular Hermitian metric $\widetilde{h}$ with semi-positive curvature, which is canonically constructed via the isomorphism ${f^{\prime}}^{*}L\sim_{{\mathbb{Q}}}K_{X^{\prime}/Y}+{\Delta}_{D^{\prime}}+M^{\prime}-\ulcorner{D^{\prime}_{-}}\urcorner$ and $b^{-1}$. The corresponding singular Hermitian metric $h_{Y}$ on $L$ (such that ${f^{\prime}}^{*}h_{Y}=\widetilde{h}$) has semi-positive curvature.
@@ 5. On finite generation problem of log canonical rings for generalized pairs
First, we see the following lemma by [BCHM].
Lemma 5.3.
Let $(X,B+M)$ be a projective generalized klt pair such that $K_{X}+B+M$ is big. Assume that $l(K_{X}+B+M)$ is Cartier divisor for an $l>0$.
Then the ring
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))$
is a finite generated $\mathbb{C}$-algebra.
Proof.
Note that by [FM, Proposition 5.1], it is enough to show the lemma for a sufficiently large and divisible $l$. Taking a log resolution, we may assume that $(X,B)$ is simple normal crossing klt pair and $M$ is nef. Since $K_{X}+B+M$ is big, we take an effective divisor $E$ and an ample divisor $A$ such that $l(K_{X}+B+M)\sim_{\mathbb{Z}}E+A$. Since $M+\frac{1}{l}A$ is ample, we can choose an effective $\mathbb{Q}$-divisor $A^{\prime}$ such that $(X,B+\frac{1}{l}(E+A^{\prime}))$ is klt and $A^{\prime}\sim_{\mathbb{Z}}lM+A$ for a sufficiently large and divisible positive integer $l$. Since
$2l(K_{X}+B+M)\sim_{\mathbb{Z}}l(K_{X}+B+\frac{1}{l}(E+A^{\prime})),$
it is enough to show the finite generation of
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+\frac{1}{l}(E+A^{\prime}))),$
which follows from [BCHM, Corollary 1.1.2]. ∎
@@
When $\kappa(K_{X}+B+M)=0$ or $\kappa(K_{X}+B_{M})<0$,
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))\simeq\mathbb{C}[t]\text{ or }\mathbb{C},$
respectively, where $\mathbb{C}[t]$ is the polynomial ring with one variable. Thus we may assume that $\kappa(X,l(K_{X}+B+M))>0$. We consider the Iitaka fibration $f:X\dashrightarrow Y$of $l(K_{X}+B+M)$ and take a higher birational model $X^{\prime}\to X$ induces a holomorphic morphism $f^{\prime}:X^{\prime}\to Y$ which is birational equivalent to $f$. Thus, by replacing, we may assume that $(X,B)$ is a simple normal crossing klt pair admitting the Iitaka fibration $f:X\to Y$ of $K_{X}+B+M$, and that $M$ has a semi-positive singular Hermitian metric with a vanishing Lelong number for the potential at every point of $X$. Now, we apply the weak semi-stable reduction theorem for $f$.
@@
We use the notation in twisted Kawamata’s semi-positivity theorem 1.1. Consider the divisor $\Delta_{Y^{\prime}}^{-}$. Since $B$ is effective, any components of $f^{\prime*}\Delta_{Y^{\prime}}^{-}$ are $\varphi$-exceptional divisors. Thus, $f^{\prime}$ is also the Iitaka fibarion of $K_{X^{\prime}}+f^{\prime*}\Delta_{Y^{\prime}}^{-}+B^{\prime}+M^{\prime}$. Therefore the finite generation of $R(K_{X}+B+M)$ is reduced to that of $R(K_{X^{\prime}}+f^{\prime*}\Delta_{Y^{\prime}}^{-}+B^{\prime}+M^{\prime}).$ And that is equivalent to the finite generation of $R(K_{Y^{\prime}}+M_{Y^{\prime}}+\Delta^{+}_{Y^{\prime}})$. Now $(Y^{\prime},\Delta_{Y^{\prime}}^{+})$ is klt and $M_{Y^{\prime}}$ is nef by Theorem 1.1. Siince $K_{Y^{\prime}}+M_{Y^{\prime}}+\Delta^{+}_{Y^{\prime}}$ is big, the desired finite generation follows from Lemma 5.3. ∎
The following is an interesting corollary for even smooth projective varieties.
@@
In the last, we discuss the finite generation of generalized canonical rings for generalized klt pairs in a general situation. Indeed we prove that for lower dimensional cases as follows:
Theorem 5.6.
Let $(X,B+M)$ be a projective generalized klt pair such that $\mathrm{dim}\,X\leq 3$. Assume that $l(K_{X}+B+M)$ is Cartier divisor for an $l>0$.
Then the ring
$\bigoplus_{m\geq 0}H^{0}(X,ml(K_{X}+B+M))$
is a finite generated $\mathbb{C}$-algebra.
Proof.
If $\kappa(X,l(K_{X}+B+M))\leq 0$, the finite generation is clear as the proof of Theorem 5.4. We may assume that $1\leq\kappa(K_{X}+B+M)\leq 2$ by Lemma 5.3. The case of $\kappa(K_{X}+B+M)=1$ is known since the finite generation is always true for a section ring of a Cartier divisor with Iitaka dimension one (cf. [W, Lemma 4.4]). When $\mathrm{dim}\,X=2$, we complete the proof by Lemma 5.3 again. Thus we may assume that $\mathrm{dim}\,X=3$ and $\kappa(K_{X}+B+M)=2$. As the proof of Theorem 5.4, by taking the Iitaka fibration and log resolution, we may assume that there exists a projective surjective morphism $f:X\to Y$ with connected fibers such that $(X,B+M)$ is generalized klt and $\mathrm{dim}\,Y=2$. For a general fiber $F$ of $f$, it holds that $\kappa((K_{X}+B+M)|_{F})=0$ since the property of Iitaka fibrations. Since $F$ is smooth projective curve, we see that $(K_{X}+B+M)|_{F}\sim_{\mathbb{Q}}0.$ Thus we can apply [FiS, Theorem 2.20].
Then there exists $B_{Y}$ and $M_{Y}$ such that $(Y,B_{Y}+M_{Y})$ is generalized klt and the finite generation for $K_{X}+B+M$ is equivalent for that for a big divisor $K_{Y}+B_{Y}+M_{Y}$ as the proof of Theorem 5.4. Since $K_{Y}+B_{Y}+M_{Y}$ is big, the desired finite generation follows from Lemma 5.3. ∎